OpenStudy (anonymous):

integrate from 1/2 to 0. arccos x dx

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+%28arccosx%29%29 then just plug it F(.5)-F(0)

OpenStudy (anonymous):

sorry i dont understand that

OpenStudy (anonymous):

\begin{eqnarray*}\int_{1/2}^0 \arccos(x)dx &=& \int_{1/2}^0 (x)'\arccos(x)dx \\ &=& x \arccos(x) \big|_{1/2}^0 + \int_{1/2}^0 \frac{x}{\sqrt{1-x^2}}dx \\ &=& -\frac{\pi}{6} -\sqrt{1-x^2}\big|_{1/2}^0 \\ &=& -\frac{\pi}{6} - 1 + \frac{\sqrt{3}}{2}.\end{eqnarray*}